# A splitting method for numerical relativistic magnetohydrodynamics

### v1

@@ 2.1 Ideal Relativistic Magnetohydrodynamics
and hence $\Pi^{\nu}_{\mbox{\tiny t}}$ is also time-like. Thus, if the numerical integration results in a space-like $\Pi^{\nu}$, the conversion of conserved quantities into primitive ones will fail. However, this is the same condition as in the numerical relativistic hydrodynamics, and hence the truncation errors arising in the numerical integration of the energy-momentum equations are unlikely to be the cause of the problems specific to the high-sigma regime of RMHD. The source of errors specific only to RMHD is the Faraday equation.
The impact of errors arising in the numerical integration of the Faraday equation on the type of the energy-momentum vector in general is complicated. Here, we investigate a special case, where its analysis is relatively simple. Namely, a state where the magnetic field vector is parallel to the velocity vector. Hence
$\Pi_{\mbox{\tiny t,0}}^{2}=\Pi_{\mbox{\tiny em,0}}^{2}+\Pi_{\mbox{\tiny pl,0}}^{2}-2{\cal E}_{\mbox{\tiny em,0}}{\cal E}_{\mbox{\tiny pl,0}}<0\,.$
Suppose that numerical integration leads to a state with the same $\Pi^{\nu}_{\mbox{\tiny t,0}}$, but the magnetic field is calculated with the relatively small error in its magnitude $\delta B>0$. Hence $\Pi_{\mbox{\tiny em}}^{2}=\Pi_{\mbox{\tiny em,0}}^{2}+\delta\Pi_{\mbox{\tiny em}}^{2}$, where $\delta\Pi_{\mbox{\tiny em}}^{2}=-B_{(0)}^{3}\delta B<0$, and ${\cal E}_{\mbox{\tiny em}}={\cal E}_{\mbox{\tiny em,0}}+\delta{\cal E}_{\mbox{\tiny em}}$, where $\delta{\cal E}_{\mbox{\tiny em}}=B_{(0)}\delta B>0$, and
$\Pi_{\mbox{\tiny t,0}}^{2}=\Pi_{\mbox{\tiny em,0}}^{2}+\delta\Pi_{\mbox{\tiny em}}^{2}+\Pi_{\mbox{\tiny pl}}^{2}-2({\cal E}_{\mbox{\tiny em,0}}+\delta{\cal E}_{\mbox{\tiny em}}){\cal E}_{\mbox{\tiny pl}}\,.$
From these expressions for the same $\Pi_{\mbox{\tiny t,0}}^{2}$, one finds
$\Pi_{\mbox{\tiny pl}}^{2}=\Pi_{\mbox{\tiny pl,0}}^{2}-2{\cal E}_{\mbox{\tiny em,0}}{\cal E}_{\mbox{\tiny pl,0}}-\delta\Pi_{\mbox{\tiny em}}^{2}+2({\cal E}_{\mbox{\tiny em,0}}+\delta{\cal E}_{\mbox{\tiny em}}){\cal E}_{\mbox{\tiny pl}}\,,$
@@
which must be negative. In this expression, $\delta{\cal E}_{\mbox{\tiny em}}\ll{\cal E}_{\mbox{\tiny em,0}}$ and can be ignored. Provided ${\cal E}_{\mbox{\tiny pl}}\ll{\cal E}_{\mbox{\tiny em}}$ the term $\Pi_{\mbox{\tiny pl,0}}^{2}\ll{\cal E}_{\mbox{\tiny em,0}}{\cal E}_{\mbox{\tiny pl}}$ and can also be ignored. These simplifications lead to the physicality condition
$2{\cal E}_{\mbox{\tiny em,0}}({\cal E}_{\mbox{\tiny pl,0}}-{\cal E}_{\mbox{\tiny pl}})>-\delta\Pi_{\mbox{\tiny em}}^{2}>0\,.$
Hence, the plasma energy has to be lower compared its original value. In turn, this condition can be satisfied only if
$2{\cal E}_{\mbox{\tiny em,0}}{\cal E}_{\mbox{\tiny pl,0}}>-\delta\Pi_{\mbox{\tiny em}}^{2}\,,$
and hence we obtain the condition on the value of the relative relative error in the magnetic field integration
$\frac{\delta B}{B_{(0)}}<\frac{{\cal E}_{\mbox{\tiny pl,0}}}{B_{(0)}^{2}}\approx\frac{\gamma^{2}}{\sigma}\,.$
Thus, the higher is the magnetisation, the more accurate the computations need to be to avoid the conversion failures.
@@ 2.2 Splitting equations of ideal RMHD into the electromagnetic field and plasma components
$\partial_{t}\left(\frac{E_{(0)}^{2}+B_{(0)}^{2}}{2}\right)+\nabla\!\cdot\!(\mn@boldsymbol{E_{(0)}}\!\times\!\mn@boldsymbol{B_{(0)}})=0\,,$ (9)
$\partial_{t}\left(\mn@boldsymbol{E_{(0)}}\!\times\!\mn@boldsymbol{B_{(0)}}\right)+\nabla\!\cdot\!\left(-\mn@boldsymbol{E}_{(0)}\otimes\mn@boldsymbol{E}_{(0)}-\mn@boldsymbol{B}_{(0)}\otimes\mn@boldsymbol{B}_{(0)}+\mn@boldsymbol{g}\left(\frac{E_{(0)}^{2}+B_{(0)}^{2}}{2}\right)\right)=0\,,$ (10)
@@
(${E_{(0)}}\!\times\!\mn@boldsymbol{B_{(1)}}+\mn@boldsymbol{E_{(1)}}\!\times\!\mn@boldsymbol{B_{(0)}}+\mn@boldsymbol{E_{(1)}}\!\times\!\mn@boldsymbol{B_{(1)}}+w\gamma^{2}\mn@boldsymbol{v}\right)=0\,,$ (15)
@@
(${E}_{(1)}\otimes\mn@boldsymbol{E}_{(0)}-\mn@boldsymbol{E}_{(0)}\otimes\mn@boldsymbol{E}_{(1)}-\mn@boldsymbol{E}_{(1)}\otimes\mn@boldsymbol{E}_{(1)}$
${B}_{(1)}\otimes\mn@boldsymbol{B}_{(0)}-\mn@boldsymbol{B}_{(0)}\otimes\mn@boldsymbol{B}_{(1)}-\mn@boldsymbol{B}_{(1)}\otimes\mn@boldsymbol{B}_{(1)}$ (16)
$\quad\quad+w\gamma^{2}\mn@boldsymbol{v}\otimes\mn@boldsymbol{v}+\mn@boldsymbol{g}\Big{[}(\mn@boldsymbol{E_{(0)}}\!\cdot\!\mn@boldsymbol{E_{(1)}})+(\mn@boldsymbol{B_{(0)}}\!\cdot\!\mn@boldsymbol{B_{(1)}})+\dfrac{E_{(1)}^{2}+B_{(1)}^{2}}{2}+p\Big{]}\Big{)}=0\,,$ (17)
$\partial_{t}\rho\gamma+\nabla\!\cdot\!\rho\gamma\mn@boldsymbol{v}=0\,,$ (18)
@@
The energy-momentum equations (15)-(17) do not involve the terms quadratic in ${B}_{0}$ and ${E}_{0}$, which are dominant in problems with high $\sigma$. As a result, the effect of the truncation error in calculations of ${B}_{0}$ on plasma parameters is reduced. The FFDE field still enters the plasma equations via linear terms. These interaction terms describe both the effect of the electromagnetic field on the plasma motion, and the effect plasma inertia on the evolution of the electromagnetic field. The two components of the electromagnetic field are also coupled via the perfect conductivity equation (20).
@@ 2.3 Numerical splitting
    $\partial_{t}\mn@boldsymbol{q}+\nabla\!\cdot\!\mn@boldsymbol{f}=0,$ (25)
@@ 2.4 Controlled energy transfer
$\delta{\cal E}_{(0)}^{n+1}>\alpha_{e}{\cal E}_{(0)}^{n+1}\,,$ (38)
where the switch-off parameter $\alpha_{e}\geq 0$. When $\alpha_{e}=0$, the transfer takes place whenever $\delta{\cal E}_{(0)}^{n+1}>0$, and when $\alpha_{e}=1$, it is turned off completely. In most of the test simulations, we used $\alpha_{e}=10^{-3}$.
@@ 3 Numerical implementation
To integrate the conservation laws of the split RMHD, we used a third-order finite-difference scheme. In this, we closely followed the scheme ECHO developed by (Del Zanna et al., 2007) for unsplit RMHD equations. There are, however, few significant differences: 1) Use of the GLM approach (Dedner et al., 2002) instead of the CT method (Evans & Hawley, 1988) to enforce the differential constraints (11) and (19); 2) Use of a novel 3rd-order WENO reconstruction algorithm; 3) Switching the DER operator (Del Zanna et al., 2007) off at shock waves to reduce numerical oscillations; 4) New variables conversion algorithm adjusted to the peculiarities of the split RMHD equations.
@@ 3.1 GLM approach
To keep the magnetic field approximately divergence-free, we follow the method called Generalised Lagrange Multiplier (GLM, Munz et al., 2000; Dedner et al., 2002). Hence, we introduce two additional dependent variables $\Phi_{(0)}$ and $\Phi_{(1)}$, one per each subsystems, and replace the Faraday equations (8,14) and the divergence-free conditions (11,19) with
$\partial_{t}\mn@boldsymbol{B}_{(s)}+\mn@boldsymbol{\nabla}\!\times\!\mn@boldsymbol{E}_{(s)}+\nabla{\Phi}_{(s)}=0\,,$ (40)
$\partial_{t}\Phi_{(s)}+\nabla\!\cdot\!\mn@boldsymbol{B}_{(s)}=-\kappa\Phi_{(s)}\,,$ (41)
In the test simulation, we use $\kappa=0.2/\Delta t$, making the $e$-folding time for $\Phi_{(s)}$ (in the case of vanishing ${B}_{(s)}$) equal to 5 integration time-steps $\Delta t$.

### v3

@@ 2.1 Ideal Relativistic Magnetohydrodynamics
Obviously, if the numerical integration results in a space-like $\Pi_{\mbox{\tiny t}}^{\nu}$, the conversion of conserved quantities into primitive ones will fail. However, this is the same as in the numerical relativistic hydrodynamics, and hence the truncation errors arising in the numerical integration of the energy-momentum equations are unlikely to be the cause of the problems specific to the high-sigma regime of RMHD. The source of errors specific only to RMHD is the Faraday equation. In the rest of this section, we demonstrate that sufficiently large errors in magnetic field can render the set of conserved variables unphysical by pushing the energy-momentum vector of plasma $\Pi_{\mbox{\tiny pl}}^{\nu}$ into the space-like domain.
We start with a physical state with the magnetic field ${B}_{0}$ and the total energy-momentum $\Pi^{\nu}_{\mbox{\tiny t,0}}$. Then we analyse other states with $\Pi_{\mbox{\tiny t}}^{\nu}=\Pi_{\mbox{\tiny t,0}}^{\nu}$ but ${B}=\mn@boldsymbol{B}_{0}+\delta\mn@boldsymbol{B}$, where the small perturbation $\delta\mn@boldsymbol{B}$ plays the role of the computational error. The aim is to determine how large this error can be before the energy-momentum vector of plasma $\Pi_{\mbox{\tiny pl}}^{\nu}$ turns space-like. In general, the impact of this error is hard to analyse. To simplify the analysis, we assume that $\delta\mn@boldsymbol{B}\parallel\mn@boldsymbol{B}_{\mbox{\tiny 0}}$ and consider only the cases where ${v}_{\mbox{\tiny 0}}$ is either parallel or perpendicular to ${B}_{\mbox{\tiny 0}}$. Since the total momentum vector
${S}_{\mbox{\tiny t}}=(B^{2}+w\gamma^{2})\mn@boldsymbol{v}-(\mn@boldsymbol{v}\!\cdot\!\mn@boldsymbol{B})\mn@boldsymbol{B}\,,$
${v}$ will remain either parallel or perpendicular to ${B}$ in the perturbed state as well. We will also assume that the magnetic field strength increases, $B=B_{\mbox{\tiny 0}}+\delta B$ where $\delta B>0$, as only this case is constraining.
When ${v}_{\mbox{\tiny 0}}\parallel\mn@boldsymbol{B}_{\mbox{\tiny 0}}$, the electromagnetic momentum ${S}_{\mbox{\tiny em,0}}=\mn@boldsymbol{0}$, and hence
$\Pi_{\mbox{\tiny t,0}}^{2}=\Pi_{\mbox{\tiny em,0}}^{2}+\Pi_{\mbox{\tiny pl,0}}^{2}-2{\cal E}_{\mbox{\tiny em,0}}{\cal E}_{\mbox{\tiny pl,0}}\,,$ (13)
and
$\Pi_{\mbox{\tiny t}}^{2}=\Pi_{\mbox{\tiny em}}^{2}+\Pi_{\mbox{\tiny pl}}^{2}-2{\cal E}_{\mbox{\tiny em}}{\cal E}_{\mbox{\tiny pl}}\,.$ (14)
Since $\Pi_{\mbox{\tiny t}}^{2}=\Pi_{\mbox{\tiny t,0}}^{2}$, this implies that
$\Pi_{\mbox{\tiny pl}}^{2}=\Pi_{\mbox{\tiny pl,0}}^{2}-\delta\Pi_{\mbox{\tiny em}}^{2}+2({\cal E}_{\mbox{\tiny em}}{\cal E}_{\mbox{\tiny pl}}-{\cal E}_{\mbox{\tiny em,0}}{\cal E}_{\mbox{\tiny pl,0}})\,,$ (15)
@@
where $-\delta\Pi_{\mbox{\tiny em}}^{2}=-(\Pi_{\mbox{\tiny em}}^{2}-\Pi_{\mbox{\tiny em,0}}^{2})=B_{0}^{3}\delta B>0$. For $\Pi_{\mbox{\tiny pl}}^{\nu}$ to be time-like, the whole expression on the right-hand side of this equation must be negative, which is the condition for the perturbed state to be physical. When ${\cal E}_{\mbox{\tiny pl}}\ll{\cal E}_{\mbox{\tiny em}}$, the term $\Pi_{\mbox{\tiny pl,0}}^{2}\ll{\cal E}_{\mbox{\tiny em,0}}{\cal E}_{\mbox{\tiny pl}}$ can be ignored. One may also ignore the small difference $\delta{\cal E}_{\mbox{\tiny em}}$ between ${\cal E}_{\mbox{\tiny em}}$ and ${\cal E}_{\mbox{\tiny em,0}}$. With these simplifications, the physicality condition reads
$2{\cal E}_{\mbox{\tiny em,0}}({\cal E}_{\mbox{\tiny pl}}-{\cal E}_{\mbox{\tiny pl,0}})-\delta\Pi_{\mbox{\tiny em}}^{2}<0\,.$ (16)
Since the last term in this expression is positive, this implies that the plasma energy in the perturbed state must be lower than the one in the original state. Moreover, this condition will not be satisfied by any ${\cal E}_{\mbox{\tiny pl}}>0$ unless
$-\delta\Pi_{\mbox{\tiny em}}^{2}<2{\cal E}_{\mbox{\tiny em,0}}{\cal E}_{\mbox{\tiny pl,0}}\,.$
Using ${\cal E}_{\mbox{\tiny em,0}}=B_{0}^{2}/2$, ${\cal E}_{\mbox{\tiny pl,0}}=w_{\mbox{\tiny 0}}\gamma_{\mbox{\tiny 0}}^{2}-p_{\mbox{\tiny 0}}\simeq w_{\mbox{\tiny 0}}\gamma_{\mbox{\tiny 0}}^{2}$, and utilising the fact that in this case $b^{2}=B^{2}$, we finally arrive to the upper limit on the maximum error in magnetic field
$\frac{\delta B}{B_{0}}\lesssim 2\frac{\gamma_{\mbox{\tiny 0}}^{2}}{\sigma_{\mbox{\tiny 0}}}\,.$ (17)
When ${v}_{\mbox{\tiny 0}}\perp\mn@boldsymbol{B}_{\mbox{\tiny 0}}$, the electromagnetic momentum does not vanish. Hence equation (15) is replaced with
$\Pi_{\mbox{\tiny pl}}^{2}$ $=\Pi_{\mbox{\tiny pl,0}}^{2}-\delta\Pi_{\mbox{\tiny em}}^{2}+$
$+2({\cal E}_{\mbox{\tiny em}}{\cal E}_{\mbox{\tiny pl}}-{\cal E}_{\mbox{\tiny em,0}}{\cal E}_{\mbox{\tiny pl,0}})+2((\mn@boldsymbol{S}_{\mbox{\tiny em,0}}\!\cdot\!\mn@boldsymbol{S}_{\mbox{\tiny pl,0}})-(\mn@boldsymbol{S}_{\mbox{\tiny em}}\!\cdot\!\mn@boldsymbol{S}_{\mbox{\tiny pl}}))\,,$ (18)
where $-\delta\Pi_{\mbox{\tiny em}}^{2}=B_{\mbox{\tiny 0}}^{3}\delta B/\gamma_{\mbox{\tiny 0}}^{4}>0$, and equation (16) with
$2{\cal E}_{\mbox{\tiny em,0}}\delta{\cal E}_{\mbox{\tiny pl}}-2(\mn@boldsymbol{S}_{\mbox{\tiny em,0}}\!\cdot\!\delta\mn@boldsymbol{S}_{\mbox{\tiny pl}})-\delta\Pi_{\mbox{\tiny em}}^{2}<0\,,$ (19)
where ${\cal E}_{\mbox{\tiny em,0}}=B_{\mbox{\tiny 0}}^{2}(1+v_{\mbox{\tiny 0}}^{2})/2$, ${S}_{\mbox{\tiny em,0}}=B_{\mbox{\tiny 0}}^{2}\mn@boldsymbol{v}_{\mbox{\tiny 0}}$, $\delta{\cal E}_{\mbox{\tiny pl}}={\cal E}_{\mbox{\tiny pl}}-{\cal E}_{\mbox{\tiny pl,0}}$, and $\delta\mn@boldsymbol{S}_{\mbox{\tiny pl}}=\mn@boldsymbol{S}_{\mbox{\tiny pl}}-\mn@boldsymbol{S}_{\mbox{\tiny pl,0}}$. When $v_{\mbox{\tiny 0}},v\ll 1$, the second term in (19) can be ignored, leading to the condition (17) with $\gamma_{\mbox{\tiny 0}}=1$. When $v_{\mbox{\tiny 0}},v\simeq 1$, one may use the approximation $\delta\mn@boldsymbol{S}_{\mbox{\tiny pl}}\simeq\delta{\cal E}_{\mbox{\tiny pl}}\mn@boldsymbol{v}_{\mbox{\tiny 0}}$, which yields $2(\mn@boldsymbol{S}_{\mbox{\tiny em,0}}\!\cdot\!\delta\mn@boldsymbol{S}_{\mbox{\tiny pl}})\simeq 2B_{\mbox{\tiny 0}}^{2}v_{\mbox{\tiny 0}}^{2}\delta{\cal E}_{\mbox{\tiny pl}}$. Substituting these into (19), one obtains the simplified physicality condition
$\frac{B_{\mbox{\tiny 0}}^{2}\delta{\cal E}_{\mbox{\tiny pl}}}{\gamma_{\mbox{\tiny 0}}^{2}}-\delta\Pi_{\mbox{\tiny em}}^{2}<0\,.$
Thus, like in the parallel case, $\delta{\cal E}_{\mbox{\tiny pl}}$ must be negative. Moreover, this condition will not be satisfied by any ${\cal E}_{\mbox{\tiny pl}}>0$ unless
$-\delta\Pi_{\mbox{\tiny em}}^{2}<\frac{B_{\mbox{\tiny 0}}^{2}{\cal E}_{\mbox{\tiny pl,0}}}{\gamma_{\mbox{\tiny 0}}^{2}}\,,$
which leads to the upper limit on the error in magnetic field
$\frac{\delta B}{B_{0}}\lesssim\frac{\gamma_{\mbox{\tiny 0}}^{2}}{\sigma_{\mbox{\tiny 0}}}\,,$ (20)
where we applied $B^{2}=\gamma^{2}b^{2}$. Interestingly, this limit differs from the one obtained for the parallel case only by the factor of two, suggesting that there is no strong dependence on the angle between the velocity and magnetic field. The accuracy constraints may be even more restrictive than those we have derived. For example, we did not take into account that the plasma energy of the perturbed state cannot fall below $D=\rho_{\mbox{\tiny 0}}\gamma_{\mbox{\tiny 0}}$, the conserved variable that has to be preserved in the perturbed state as well.
In the standard approach, one may try to tackle the issue of conversion failures via increasing the accuracy of the magnetic field integration. In the splitting approach, we seek to reduce the impact of the errors in the magnetic field on the energy-momentum of plasma.
@@ 2.2 Splitting equations of ideal RMHD into the electromagnetic field and plasma components
$\partial_{t}\left(\frac{E_{(0)}^{2}+B_{(0)}^{2}}{2}\right)+\mn@boldsymbol{\nabla}\!\cdot\!(\mn@boldsymbol{E}_{(0)}\!\times\!\mn@boldsymbol{B}_{(0)})=0\,,$ (22)
$\partial_{t}\left(\mn@boldsymbol{E_{(0)}}\!\times\!\mn@boldsymbol{B_{(0)}}\right)+$
+(${E}_{(0)}\otimes\mn@boldsymbol{E}_{(0)}-\mn@boldsymbol{B}_{(0)}\otimes\mn@boldsymbol{B}_{(0)}+\mn@boldsymbol{g}\left(\frac{E_{(0)}^{2}+B_{(0)}^{2}}{2}\right)\right)=0\,,$ (23)
@@
+(${E}_{(0)}\!\times\!\mn@boldsymbol{B}_{(1)}+\mn@boldsymbol{E}_{(1)}\!\times\!\mn@boldsymbol{B}_{(0)}+\mn@boldsymbol{E}_{(1)}\!\times\!\mn@boldsymbol{B}_{(1)}+w\gamma^{2}\mn@boldsymbol{v}\right)=0\,,$ (28)
@@
+(${E}_{(1)}\otimes\mn@boldsymbol{E}_{(0)}-\mn@boldsymbol{E}_{(0)}\otimes\mn@boldsymbol{E}_{(1)}-\mn@boldsymbol{E}_{(1)}\otimes\mn@boldsymbol{E}_{(1)}$
${B}_{(1)}\otimes\mn@boldsymbol{B}_{(0)}-\mn@boldsymbol{B}_{(0)}\otimes\mn@boldsymbol{B}_{(1)}-\mn@boldsymbol{B}_{(1)}\otimes\mn@boldsymbol{B}_{(1)}$
+${g}\Big{[}\mn@boldsymbol{E}_{(0)}\!\cdot\!\mn@boldsymbol{E}_{(1)}+\mn@boldsymbol{B}_{(0)}\!\cdot\!\mn@boldsymbol{B}_{(1)}+\dfrac{E_{(1)}^{2}+B_{(1)}^{2}}{2}\Big{]}$
$\quad\quad\,\,\,+w\gamma^{2}\mn@boldsymbol{v}\otimes\mn@boldsymbol{v}+\mn@boldsymbol{g}p\Big{)}=0\,,$ (29)
$\partial_{t}\rho\gamma+\mn@boldsymbol{\nabla}\!\cdot\!\rho\gamma\mn@boldsymbol{v}=0\,,$ (31)
@@
The energy-momentum equations (28)-(LABEL:eq:mom1) do not involve the terms quadratic in ${B}_{(0)}$ and ${E}_{(0)}$, which are dominant in problems with high $\sigma$. As a result, the effect of the truncation error in calculations of ${B}_{(0)}$ on plasma parameters is reduced. The FFDE field still enters the plasma equations via linear terms. These interaction terms describe both the effect of the electromagnetic field on the plasma motion, and the effect plasma inertia on the evolution of the electromagnetic field. The two components of the electromagnetic field are also coupled via the perfect conductivity equation (33).
@@ 2.3 Numerical splitting
    $\partial_{t}\mn@boldsymbol{q}+\mn@boldsymbol{\nabla}\!\cdot\!\mn@boldsymbol{f}=0,$ (38)
@@ 2.4 Controlled energy transfer
$\delta{\cal E}_{(0)}^{n+1}>\alpha_{\mbox{\tiny e}}{\cal E}_{(0)}^{n+1}\,,$ (51)
where the switch-off parameter $\alpha_{\mbox{\tiny e}}\geq 0$. When $\alpha_{\mbox{\tiny e}}=0$, the transfer takes place whenever $\delta{\cal E}_{(0)}^{n+1}>0$, and when $\alpha_{\mbox{\tiny e}}=1$, it is turned off completely. In most of the test simulations, we used $\alpha_{\mbox{\tiny e}}=10^{-3}$.
@@ 3 Numerical implementation
To integrate the conservation laws of the split RMHD, we used a third-order finite-difference scheme. In this, we closely followed the scheme ECHO developed by (Del Zanna et al., 2007) for unsplit RMHD equations. There are, however, few significant differences: 1) Use of the GLM approach instead of the CT method to enforce the differential constraints (24) and (32). GLM delivers similar accuracy to CT (Mignone et al., 2010) for ideal MHD, but it is much easier to implement; 2) Use of a novel 3rd-order WENO reconstruction algorithm; 3) Switching the DER operator (Del Zanna et al., 2007) off at shock waves to reduce numerical oscillations; 4) New variables conversion algorithm adjusted to the peculiarities of the split RMHD equations.
@@ 3.1 GLM approach
To keep the magnetic field approximately divergence-free, we follow the method called Generalised Lagrange Multiplier (GLM, Munz et al., 2000; Dedner et al., 2002). Hence, we introduce two additional dependent variables $\Phi_{(0)}$ and $\Phi_{(1)}$, one per each subsystems, and replace the Faraday equations (21,27) and the divergence-free conditions (24,32) with
$\partial_{t}\mn@boldsymbol{B}_{\mbox{\tiny(s)}}+\mn@boldsymbol{\nabla}\!\times\!\mn@boldsymbol{E}_{\mbox{\tiny(s)}}+\nabla{\Phi}_{\mbox{\tiny(s)}}=0\,,$ (53)
$\partial_{t}\Phi_{\mbox{\tiny(s)}}+\mn@boldsymbol{\nabla}\!\cdot\!\mn@boldsymbol{B}_{\mbox{\tiny(s)}}=-\kappa\Phi_{\mbox{\tiny(s)}}\,,$ (54)
where $s=0,1$. In the test simulation, we use $\kappa=0.2/\Delta t$, making the $e$-folding time for $\Phi_{\mbox{\tiny(s)}}$ (in the case of vanishing ${B}_{\mbox{\tiny(s)}}$) equal to 5 integration time-steps $\Delta t$.